# Multi-Agent Inverse Reinforcement Learning in Real World Unstructured Pedestrian Crowds

### v2

@@ IV Multi-agent MaxEnt IRL for Unstructued Pedestrian Crowds
In this section, we discuss the technical details of our proposed multi-agent MaxEnt IRL approach. We visually summarize the algorithm in Figure 2. In MaxEnt IRL [26], the goal is to learn a cost function that applies either to a single agent or is designed to collectively model a group of individuals. However, individuals in groups behave differently according to their nature. For instance, some individuals might be more risk averse than others, and therefore, are more likely to have a higher reward for collision avoidance. Others might place greater importance on reaching their goals faster and place greater weight for the cost of distance to goal. Therefore, we propose the following multi-agent formulation of MaxEnt IRL for unstructured pedestrian crowds.
@@
However, the constant velocity dynamics assumption, combined with the fact that humans typically walk along linear paths at near identical speeds on average, results in a linear approximation of the cost function in Equation (16). Under constant velocity dynamics and linear motion, the state evolution is linear. The deviations $\delta_{x_{t}}$ from the trajectory $\bar{x}_{t}$ remain small and linear. Consequently, the second-order derivatives (i.e., the entries of $H_{t}^{i}$) become very small since there is minimal curvature in the cost function along linear trajectories. Mathematically, the curvature of the cost function is captured by the eigenvalues of the Hessian $H_{t}^{i}$. If the cost function has little curvature (as it does under linear motion), these eigenvalues are close to zero, resulting in a non positive semi-definite (PSD) matrix.
@@
where $g\left(\textrm{diag}(\Sigma^{i}_{t})\right)$ produces a vector that adjusts the diagonal elements of $\Sigma^{i}_{t}$. This operation ensures that $\widetilde{\Sigma}^{i}_{t}$ is a positive semi-definite (PSD) matrix, thus preserving the stochastic nature of the policy while making the problem more tractable. The diagonal elements of $\Sigma^{i}_{t}$, denoted as $\sigma^{ij},i=j$, quantify the uncertainty or randomness in the action that agent $i$ takes toward agent $j$. By applying the function $g$ to these diagonal elements, we effectively increase the variance (uncertainty) of the agent’s actions. This modification shifts towards a more tractable model at the cost of potentially capturing a less accurate representation of agent behaviors, as seen in our evaluation.
V Evaluation
We compare our multi-agent MaxEnt IRL approach with MaxEnt IRL, behavior cloning, and a generative modeling approach and aim to understand how our method compares to these baselines in real world pedestrian crowds.
V-A Data Pre-processing
Dataset Description
Structured Unstructured Agents involved Density (# agents / sq. meter)
ETH/UCY [70] - -
SCAND [6] - -
JRDB [71] - -
Waymo [25] ✓(0.336 bits) Vehicles $1$
INTERACTION [72] ✓(0.475 bits) Vehicles $1$
Speedway ✓(1.061 bits) Pedestrians and Vehicles $\mathbf{3-4}$
TABLE I: Comparing our dataset with state-of-the-art motion datasets: Our dataset features trajectories of pedestrians as well as vehicles characterized by higher entropy (greater unpredictability) and increased density.
Figure 3: Cumulative RMSE distribution demonstrates the percentage of trajectories below a certain RMSE threshold; steeper curves indicate a more effective baseline.
(a) Multi-agent Max-Ent IRL
(b) Max-Ent IRL [27]
(c) Behavior cloning [28]
(d) PECNet [30]
[…]
We collected an unstructured real world pedestrian crowd dataset at one of the busiest intersections on The University of Texas at Austin campus using a BlueCity [^2] https://ouster.com/products/software/blue-city Velodyne LiDAR mounted at the intersection. Our dataset consists of more tightly coupled agent interactions compared to existing motion planning datasets [6, 71, 25]. Specifically, in Figure LABEL:fig:_lines, we visualize the difference between our dataset and an intersection chosen at random from the Waymo dataset, which is considered one of state-of-the-art benchmarks for motion prediction [25]. We observe that trajectories in our dataset are denser and more unstructured (higher entropy), making it an ideal benchmark to test the ability of algorithms to handle unstructured crowds.
V-B Baselines and Evaluation Metric
We compare our approach, which we call MAIRL (multi-agent IRL), with the following baselines [^3] Note that these baselines were selected by identifying state-of-the-art methods that $(i)$ were compatible with our data format and $(ii)$ had open source implementations.: single-agent MaxEnt IRL [26], a state-of-the-art behavior cloning (trajectory prediction) algorithm [28], and a generative modeling algorithm (PECNet [30]). We directly use the opensource software for all these baselines and train them on our dataset to present a direct evaluation. We set the learning rate to $3\times 10^{-4}$ and perform mini-batch training with a batch size of $64$. A $60-40$ cross-validation split is used for evaluation. To prevent overfitting, we apply a dropout rate of $0.2$. Training utilizes the AdamW optimizer with the Mean Squared Error (MSE) loss function. All training was performed on an NVIDIA GeForce RTX 2080 GPU machine and the average time to train the baselines was in the order of minutes. Since we do not know the true cost functions of the pedestrians, we are unable to measure some important social navigation metrics such as “comfort”, although we acknowledge those are important as well, and we leave that for future work. Instead, we directly minimize the root mean square error between the demonstration trajectories and the trajectories generated via the policy $\mathcal{F}^{i}_{\theta^{i}}$. This is a standard evaluation protocol followed in the navigation literature [26, 29].
V-C Results and Discussion
@@
TABLE II: Quantitative results of our proposed method compared to state-of-the-art baselines on five benchmark datasets. All numbers are in meters.
@@
JRDB [71] $3.939$ $2.671$ $3.498$ $2.646$ $4.249$ $2.558$ $2.506$ $2.640$
SCAND [6] $0.55/1.20$ $0.50/1.10$ $0.44/0.90$ $0.36/0.81$ $0.42/0.85$ $0.23/0.67$
TABLE III: EFE results on the JRDB and SCAND dataset.
V-C1 Results
In Figure 1, we plot the root mean square error (RMSE) between the planned and ground truth trajectories as a function of the timesteps. Here, a timestep refers to the LiDAR’s capture frequency, and not in terms of seconds. We observe that MAIRL incurs the lowest RMSE compared to behavior cloning, MaxEnt IRL, and PECNet, which implies that MAIRL most accurately recovers the ground truth reward functions of multiple agents. In Figure 3, we plot a cumulative distribution frequency (CDF) that measures the percentage of trajectories that fall below a certain RMSE threshold; steeper curves imply a more effective algorithm as it indicates that that the algorithm was able to generate more low-RMSE trajectories. The CDF plot is useful because the RMSE curve alone may not convey the full picture. For example, a baseline could yield a poor RMSE, and be declared as weak algorithm, but a low RMSE could result from one or two outlier trajectories with an particularly high RMSE, thereby skewing the average against its favor.
@@ V-C1 Results
V-C2 Discussion
Our findings demonstrate that multi-agent MaxEnt IRL outperforms MaxEnt IRL, particularly when dealing with diverse trajectories from multiple agents. Due to the single-agent nature of MaxEnt IRL, it aligns the predicted trajectories with the ground truth demonstrations for at most $1$ agent (the blue agent), and fails for the green (extreme deviation) and the red (not reached goal) agents (Figure 4(b)). Behavior cloning encounters substantial difficulties, on account of the inherently non-linear pedestrian trajectories. Note that this is in contrast to many successful results in the autonomous driving domain, where agents follow structured motion along straight lanes and roads. Furthermore, our findings suggest there exists a notable exception when the models are constrained to predict solely the velocity, provided that either the heading angle or an endpoint is supplemented as ground truth to the training model (as is in the case of PECNet). We omit the results here due to lack of space, but we observed that behavior cloning performs better, when predicting only the velocity, although we acknowledge that providing the heading angle as ground truth is impractical in real world scenarios. Finally, we note that even though MAIRL outperforms all three baselines in terms of aligning with the ground truth demonstrations, it still remains imperfect in its own alignment due to the reduction in accuracy resulting from the Tractability-Rationality Trade-off, as discussed in Section III.
@@ VI Conclusion and Open Questions
We proposed a novel approach for multi-agent MaxEnt inverse reinforcement learning for real world unstructured pedestrian crowds. Our key insight was a mathematical trick that enabled tractability of multi-agent MaxEnt IRL at the cost of a slight reduction in accuracy. We used a dynamic game theoretic solver to compute optimal policies that maximize the reward functions learned via our approach. We evaluated this new approach using a dataset collected at a busy intersection on a University campus and showed that despite the accuracy trade-off, our approach outperforms state-of-the-art imitation learning baselines such as single-agent MaxEnt IRL, behavior cloning, and generative modeling. Our findings highlighted two key insights: $(i)$ our approach yields the lowest error between the policy-generated trajectories and the ground truth trajectories, $(ii)$ our approach has the highest percentage of low-RMSE trajectories. Future work includes finding ways to solve Equation (17) in unstructured pedestrian crowds without relaxing bounded rationality. Another open problem is to compute an exact bound for the tractability-rationality tradeoff trick (Section IV). In addition, incorporating external factors like social contexts and the environment in modeling pedestrian behaviors can yield more holistic and precise predictions, providing a more comprehensive view of motion dynamics in various environments.

### v3

@@ IV Multi-agent MaxEnt IRL for Unstructued Pedestrian Crowds
We visually summarize the algorithm in Figure 1. This section begins with a brief motivation for Multi-agent MaxEnt IRL, followed by describing the algorithm in detail, which includes the Tractability-Rationality trade-off trick which is depicted in the yellow box in Figure 1.
In MaxEnt IRL [26], the goal is to learn a cost function that applies either to a single agent or is designed to collectively model a group of individuals. However, individuals in groups behave differently according to their nature. For instance, some individuals might be more risk averse than others, and therefore, are more likely to have a higher reward for collision avoidance. Others might place greater importance on reaching their goals faster and place greater weight for the cost of distance to goal. In [36], Mehr et al. introduced the following multi-agent formulation of MaxEnt IRL for unstructured pedestrian crowds.
@@
Our assumption of constant velocity dynamics in Equation (14), combined with the fact that humans typically walk along linear paths at near identical speeds on average, results in a linear approximation of the cost function in Equation (15). Under constant velocity dynamics and linear motion, the state evolution is linear. The deviations $\delta_{x_{t}}$ from the trajectory $\bar{x}_{t}$ remain small and linear. Consequently, the second-order derivatives i.e., the entries of $\left(H_{t}^{i}\right)$, become very small since there is minimal curvature in the cost function along linear trajectories. Mathematically, the curvature of the cost function is captured by the eigenvalues of the Hessian $H_{t}^{i}$. If the cost function has little curvature (as it does under linear motion), these eigenvalues are close to zero, resulting in a non positive semi-definite (PSD) matrix.
@@
TABLE I: Quantitative results (ADE/FDE) of our proposed method compared to state-of-the-art baselines on five benchmark datasets. All results are in meters. Lower is better
where $g\left(\textrm{diag}(\Sigma^{i}_{t})\right)$ produces a vector that adjusts the diagonal elements of $\Sigma^{i}_{t}$. This operation ensures that $\widetilde{\Sigma}^{i}_{t}$ is a positive semi-definite (PSD) matrix, thus preserving the stochastic nature of the policy while making the problem more tractable. The diagonal elements of $\Sigma^{i}_{t}$, denoted as $\sigma^{ij},i=j$, quantify the uncertainty or randomness in the action that agent $i$ takes toward agent $j$. By applying the function $g$ to these diagonal elements, we essentially decrease the variance (uncertainty) of the agent’s actions. This modification shifts towards a more tractable model at the cost of potentially capturing a less accurate representation of agent behaviors, as seen in our evaluation.
@@
JRDB [83] $3.939$ $2.671$ $3.498$ $2.646$ $4.249$ $2.558$ $\mathbf{2.506}$ $2.640$
SCAND [6] $0.55/1.20$ $0.50/1.10$ $0.44/0.90$ $0.36/0.81$ $0.42/0.85$ $\mathbf{0.23/0.67}$
TABLE II: EFE (End-to-end Forecasting Error) results on the JRDB and SCAND dataset. All results are in meters/ Lower is better.
V Evaluation
(a) Our dataset
(b) Waymo/INTERACTION
Figure 2: Visualizing and comparing the entropy of trajectories in the Speedway dataset ($1.061$ bits) with the entropy of trajectories in the Waymo and INTERACTION datasets ($0.336$ and $0.475$ bits, respectively). We observe that trajectories in the Speedway dataset are denser, more unstructured, and have a higher entropy.
We compare our multi-agent MaxEnt IRL approach with single-agent MaxEnt IRL and state-of-the-art trajectory planning and prediction approaches and aim to understand how our method compares to these baselines on popular autonomous driving and pedestrian crowd datasets.
V-A Data Pre-processing
We collected an unstructured real world pedestrian crowd dataset at one of the busiest intersections on a University campus using a BlueCity [^2] https://ouster.com/products/software/blue-city Velodyne LiDAR mounted at the intersection. Our dataset consists of more tightly coupled agent interactions compared to existing motion planning datasets [6, 83, 25].
V-B Baselines and Evaluation Metric
We compare our approach, which we call MAIRL (multi-agent IRL), with state-of-the-art trajectory prediction methods that have consistently performed well on leading benchmarks including the JRDB and ETH/UCY. To compare MAIRL with these methods [^3] Note that for comparison on the Speedway dataset, baselines were selected by identifying state-of-the-art methods that $(i)$ were compatible with our input data format and $(ii)$ had open source implementations. on the Speedway dataset, we directly use the opensource software for these baselines and train them on our dataset to present a direct evaluation. We set the learning rate to $3\times 10^{-4}$ and perform mini-batch training with a batch size of $64$. A $60-40$ cross-validation split is used for evaluation. To prevent overfitting, we apply a dropout rate of $0.2$. Training utilizes the AdamW optimizer with the Mean Squared Error (MSE) loss function. All training was performed on an NVIDIA GeForce RTX 2080 GPU machine and the average time to train the baselines was in the order of minutes.
Following standard evaluation procedures followed by the literature [26, 36], we measure the average displacement error (ADE) and the final displacement error (FDE) between the demonstration trajectories and the trajectories generated via the policy $\mathcal{F}^{i}_{\theta^{i}}$.
(a) Multi-agent Max-Ent IRL
(b) Max-Ent IRL [27]
@@ V-B Baselines and Evaluation Metric
Figure 4: Cumulative RMSE distribution demonstrates the percentage of trajectories below a certain RMSE threshold; steeper curves indicate a more effective learner.
V-C Results and Discussion
Table I compares (in order of better performance) the ADE/FDE (lower is better) of MAIRL with state-of-the-art trajectory prediction and planning methods on the well-known ETH/UCY datasets as well as on Speedway, in meters. The horizontal rule in the middle of the table divides the methods between ones that could be tested on the Speedway dataset versus ones that could not3. The results show that our MAIRL outperforms the top methods on the Speedway dataset, and is comparable on the ETH/UCY datasets, only performing worse than Expert-Goals [24] and Trajectron++ [23], two well-adopted methods for deep learning-based trajectory prediction. Table II additionally compares MAIRL on the JRDB and SCAND datasets based on the End-to-End Forecasting Error (EFE), a specific metric used by the JRDB benchmakr, but is similar to the ADE/FDE metrics, again, in meters. MAIRL performs the best on SCAND, and comes in $3^{\textrm{rd}}$ on JRDB, behind Social-Pose [82] and HST [22], two methods that additionally take into account the skeleton pose of the pedestrians. In both Tables I and II, note that we also compare MAIRL with single agent MaxEnt IRL, which performs considerably worse.
[…]
In Figure 4, we plot a cumulative distribution frequency (CDF) that measures the percentage of trajectories that fall below a certain error threshold; steeper curves imply a more effective algorithm as it indicates that that the algorithm was able to generate more low-ADE trajectories. The CDF plot is useful because the ADE/FDE metrics alone may not convey the full picture. For example, a baseline could yield a poor ADE, and be declared as weak algorithm, but a low ADE could result from one or two outlier trajectories with an particularly high ADE, thereby skewing the average against its favor.
@@ VI Conclusion and Open Questions
We show that multi-agent MaxEnt inverse reinforcement learning provides accurate predictions of pedestrians’ motion in real world unstructured pedestrian crowds with only few training trajectories. Our key insight was a mathematical trick that enabled tractability of multi-agent MaxEnt IRL at the cost of a slight reduction in accuracy. We used a dynamic game theoretic solver to compute optimal policies that maximize the reward functions learned via our approach. We evaluated this new approach using a dataset collected at a busy intersection on a University campus and showed that despite the accuracy trade-off, our approach outperforms state-of-the-art imitation learning baselines such as single-agent MaxEnt IRL, behavior cloning, and generative modeling. Our findings highlighted that our approach yields the lowest error between the policy-generated trajectories. Future work includes finding ways to solve Equation (16) in unstructured pedestrian crowds without relaxing bounded rationality. Another open problem is to compute an exact bound for the tractability-rationality tradeoff trick (Section IV). In addition, incorporating external factors like social contexts and the environment in modeling pedestrian behaviors can yield more holistic and precise predictions, providing a more comprehensive view of motion dynamics in various environments.